The perimeter of a playing field for a certain sport is
244 ft. The field is a​ rectangle, and the length is
46 ft longer than the width. Find the dimensions.

Answer by math_helper(2461)


2L+2W = 244 ==> L+W = 122 (1)
L = W+46 (2)
Substitute for L from (2) into (1):
W+46+W = 122
2W+46 = 122
...
Solve for W=38ft and L=84ft